# How Much are LLMs Contaminated?

### v1

@@ Abstract
With the rise of Large Language Models (LLMs) in recent years, new opportunities are emerging, but also new challenges, and contamination is quickly becoming critical. Business applications and fundraising in AI have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a critical issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes the entire progress in the field of NLP, yet, there remains a lack of methods on how to efficiently address contamination, or a clear consensus on prevention, mitigation and classification of contamination. In this paper, we survey all recent work on contamination with LLMs, and help the community track contamination levels of LLMs by releasing an open-source Python library named LLMSanitize implementing major contamination detection algorithms, which link is: https://github.com/ntunlp/LLMSanitize.
@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), Large Language Models (LLMs) have emerged as pivotal tools, driving innovation across a wide spectrum of applications, from natural language processing and automated content creation (Achiam et al., 2023; Ding et al., 2023; Qin et al., 2023; Betker et al., 2023; Ding et al., 2024) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023a; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored output (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination—be it intentional manipulation or inadvertent inclusion of biased, inaccurate, or misleading information—can significantly impair the models’ effectiveness and reliability.
@@
Contamination is a broad issue that we divide into data contamination and model contamination. On one hand, data contamination refers to the scenario where an evaluation dataset is (partly) included in the training set, resulting in evaluation results not properly assessing the performance of LLMs. On the other hand, model contamination stands for the phenomenon where the model has seen a specific evaluation dataset during pre-training or fine-tuning, leading to biased results on this dataset. Additionally, if we consider whether the contaminated data contains only inputs or both inputs and labels, contamination includes input contamination and input+label contamination. Depending on the level of access to the model, model contamination detection mainly focuses on white-box detection where we have full access to the model, e.g., a local model, and black-box detection where only the API access is available, e.g., GPT-4 (Achiam et al., 2023).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Contamination, parent [Data Contamination, fill=yellow!30, domain_box [String Matching, fill=orange!30, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), models ] ] [Embeddings Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), models ] ] [LLM-based Data Contamination Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023b), Clean-Eval (Zhu et al., 2023), models] ] ] [Model Contamination, fill=yellow!30, domain_box [Performance Analysis, fill=orange!30, domain_box [Task contamination (Li & Flanigan, 2023), Pre-training GPT-2 on contaminated data (Jiang et al., 2024b), models ] ] [Model Completion, fill=orange!30, domain_box [Model Memorization, fill=red!20, domain_box [Injecting test set in training set (Magar & Schwartz, 2022), Completion prompting (Anil et al., 2023), models ] ] [Prompt Engineering, fill=red!20, domain_box [Repetition attack (Nasr et al., 2023), According to prompting (Weller et al., 2023), models ] ] [Model Likelihood, fill=red!20, domain_box [Min-KProb (Shi et al., 2024), Sharded likelihood (Oren et al., 2023), models ] ] ] [LLM-based Model Contamination Detection, fill=orange!30, domain_box [Guided prompting (Golchin & Surdeanu, 2023b), Data Contamination Quiz (Golchin & Surdeanu, 2023a), models] ] ] ]
Figure 1: Classification of the different types of contamination detection methods, with (non-exhaustive lists) examples of methods within each category.
@@
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In contexts where precision and trustworthiness are non-negotiable, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. The credibility of LLM outputs can dramatically influence investor confidence, shaping perceptions of innovation, technological prowess, and, ultimately, financial prospects. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination.
@@
This landscape necessitates a comprehensive survey and a shared integrated resource of contamination detection in LLMs. We delineate the scope and nature of contamination challenges, identifying their sources, types, and impacts on model performance and applicability. In doing so, we not only contribute to the academic and practical understanding of data contamination issues but also highlight the pressing need for strategies that mitigate these risks, ensuring that the deployment of LLMs aligns with principles of integrity, fairness, and economic sustainability.
@@
  • We categorize contamination into two broad types of use cases, namely data and model contamination, and review all existing work in each category. We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
  • We introduce LLMSanitize, an open-source Python library aimed at helping the community centralize a shared implementation of all data and model contamination methods.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to data contamination. Section 3 addresses the issue of model contamination, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices. Section 5 introduces our library LLMSanitize implementing contamination detection algorithms.
2 Data Contamination
In data contamination detection, given a training dataset $D$ and an evaluation dataset $D_{E}$, the objective is to check whether $D\cap D_{E}$ is empty. In general, checking data contamination is primordial to prevent memorization of contaminated data, which could skew performance on the evaluation benchmarks.
@@ 2.1 String Matching
Several LLM training reports assess the level of contamination between their downstream evaluation datasets and their pre-training set through string-matching, the most straightforward data contamination assessment approach. Here we review all such methods.
GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. As a result, they discover 1-6% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training, with an average of 3.2%. This method also exposes contamination between test sets and training sets of these datasets.
GPT-3 (Brown et al., 2020) defines a clean version of each evaluation dataset by removing all data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). A conservative overlap analysis discovers large contamination problems for common datasets, showing some Wikipedia language modeling benchmarks, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) to be nearly entirely contaminated.
Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%.
PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Like GPT-3, they conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees.
Since then, GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data with substring match: a match is identified if any of three 50-characters randomly sampled substrings from the evaluation data point (or the entire data point if it smaller than this) is a substring of the training set. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results.
Llama-2 (Touvron et al., 2023) refines the methods above by using tokens instead of words to measure contamination, and establishes that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. From there, the contamination level of a sample is defined as its percentage of contaminated tokens, yielding a granular definition of data contamination.
Li (2023b) computes the METEOR (Banerjee & Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and considers those with scores over 0.75 as contaminated. Varying contamination levels are identified, ranging from 1% to 47% across popular benchmarks. Similar to PaLM and GPT-4, the study also finds that data contamination does not necessarily lead to increased performance metrics.
Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then split retrieved documents into 13-gram chunks, and compute overlap metrics between these chunks and the evaluation data point’s chunks. Results show that TruthfulQA exhibits high overlap with the pre-training datasets.
@@ 2.2 Embeddings Similarity
Beyond simple string matching, which fails on paraphrased samples, computing cosine similarity between embeddings offers an attractive alternative, as this process is more robust to vocabulary changes, including paraphrases or synonyms.
Lee et al. (2023) prevent contamination in their newly introduced Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers & Gurevych, 2019)) greater than 80% against any training item.
Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their fine-tuning set CodeExercises and the evaluation set HumanEval (Chen et al., 2021). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from (Riddell et al., 2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). They show that widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 LLM-based data contamination detection
String-matching could struggle to detect synthetic training data produced by LLMs. It is also challenging to choose an appropriate similarity threshold in semantic similarity search approaches using embeddings. To improve robustness, some methods attempt to use LLMs themselves to detect contamination, leveraging their high performance at paraphrasing and paraphrase detection (Witteveen & Andrews, 2019; Abaskohi et al., 2023).
@@ 2.3 LLM-based data contamination detection
Zhu et al. (2023) introduce Clean-Eval, a benchmark to assess the decontaminated performance of LLMs. Data points from popular benchmarks (e.g., MMLU) are paraphrased with GPT-3.5, and the LLM is evaluated on the paraphrased samples with the lowest BLEURT score with regards to the original ground truth. Results show a drop in performance for GPT-3.5 and Llama-2 on the Clean-Eval data, confirming prior contamination.
3 Model Contamination
In model contamination, a model $M$ trained on a training set $D_{M}$ is applied to an evaluation dataset $D_{E}$ and the task is to determine whether $D_{M}\cap D_{E}$ is empty. This exposure to evaluation data during training raises concerns about the model’s ability to generalize to unseen data, as its performance could be artificially inflated. Compared to data contamination, the training set is dependent on the model, and could be unknown, as with GPT-3.5 and GPT-4.
@@ 3.1 Performance Analysis
We first review a line of research assessing model contamination through performance analysis: $M$ is applied to evaluation datasets of different timestamps, including recent datasets which did not exist at pre-training time, forfeiting any contamination. A drop in performance on these recent data points suggests previous contamination.
Roberts et al. (2023) introduce a comparative analysis of the efficacy in code generation tasks between GPT-4 and GPT-3.5-Turbo, utilizing datasets from periods before and after a specified cut-off date. Through a meticulous examination of empirical evidence, the study reveals a compelling correlation between the likelihood of exposure to coding problem platforms such as Codeforces and Project Euler by these LLMs and the observed test case pass rates. Specifically, the analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks.
Li & Flanigan (2023) present four innovative methods to detect task contamination in LLMs which are crucial for maintaining the integrity of machine learning processes. These methods include (i) training data inspection to identify direct task examples within the data, (ii) task example extraction from instruction-tuned models, (iii) membership inference to detect exact content matches suggesting data inclusion, and (iv) chronological analysis to assess contamination through performance metrics against known dataset timelines. This multifaceted approach provides a comprehensive toolkit for researchers and developers to safeguard against inadvertent task leakage, thereby enhancing LLMs’ reliability.
Jiang et al. (2024b) conduct a comprehensive evaluation of data contamination in language model pre-training through experiments on various public datasets. The research highlights the critical issue of evaluation data leakage and its influence on language model performance, utilizing a novel approach by pre-training models from scratch. Findings reveal a U-shaped relationship between the extent of data contamination and model performance, challenging existing n-gram based definitions of contamination as inadequate. This underscores the necessity for more robust contamination detection methods.
3.2 Model Completion
An intuitive approach to detect model contamination in LLMs is to carefully analyze their outputs in controlled prompting setups. Outputs on contaminated data points may be too close or exactly match training data, or be subject to very high model confidence.
3.2.1 Model Memorization

### v3

# How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library

@@ Abstract
With the rise of Large Language Models (LLMs) in recent years, abundant new opportunities are emerging, but also new challenges, among which contamination is quickly becoming critical. Business applications and fundraising in AI have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a major issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes the entire progress in the field of NLP, yet, there remains a lack of methods on how to efficiently detect contamination. In this paper, we survey all recent work on contamination detection with LLMs, and help the community track contamination levels of LLMs by releasing an open-source Python library named LLMSanitize implementing major contamination detection algorithms, available here: https://github.com/ntunlp/LLMSanitize
How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library
Mathieu Ravaut1,2, Bosheng Ding1, Fangkai Jiao1,2, Hailin Chen1,3, Xingxuan Li1, Ruochen Zhao1, Chengwei Qin1, Caiming Xiong3, Shafiq Joty1,3† 1 Nanyang Technological University, Singapore 2 Institute of Infocomm Research (I2R), ASTAR, Singapore 3 Salesforce Research {mathieuj001,bosheng001,fangkai002,hailin001,xingxuan001,ruochen002,chengwei003}@e.ntu.edu.sg {cxiong,sjoty}@salesforce.com
@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), Large Language Models (LLMs) have emerged as pivotal tools, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023a; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored output (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness and reliability.
@@
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points. Additionally, we highlight the following orthogonal dimension to the duality explained above: if we consider whether the contaminated data contains only inputs or both inputs and labels, contamination includes input contamination and input+label contamination. A third dimension to the problem is the level of access to the LLM itself in closed-data setups: white-box detection where we have full access to the model, e.g., a local model, and black-box detection where only the API access is available, e.g., GPT-4 (Achiam et al., 2023).
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Contamination
Detection Methods, parent [Open-Data
Contamination Detection, fill=yellow!30, domain_box [String
Matching, fill=orange!30, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), PaLM (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023), Dodge et al. (2021); Li (2023b); Deng et al. (2023), models ] ] [Embeddings
Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), Riddell et al. (2024), models ] ] [Model-based Open-
Data Contamination Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023b), models] ] ] [Closed-Data
Contamination Detection, fill=yellow!30, domain_box [Performance
Analysis, fill=orange!30, domain_box [Roberts et al. (2023); Li and Flanigan (2023); Jiang et al. (2024b), models ] ] [Model
Memorization, fill=orange!30, domain_box [TS-Guessing (Deng et al., 2023), Elangovan et al. (2021); Magar and Schwartz (2022); Gemma Team Google DeepMind (2024); Nasr et al. (2023); Weller et al. (2023); Chang et al. (2023); Ranaldi et al. (2024), models ] ] [Model
Confidence, fill=orange!30, domain_box [Min-KProb (Shi et al., 2024), CDD and TED (Dong et al., 2024), Li (2023a); Oren et al. (2023); Xu et al. (2024b), models ] ] [Model-based Closed-
Data Contamination Detection, fill=orange!30, domain_box [Guided prompting (Golchin and Surdeanu, 2023b), Tu et al. (2024), models] ] ] ]
Figure 1: Classification of the different types of contamination detection methods.
@@
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In high-stakes scenarios where imprecision can have dramatic consequences, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination.
@@
This landscape necessitates a comprehensive survey and a shared integrated resource of contamination detection in LLMs. We delineate the scope and nature of contamination detection challenges, identifying their types. In doing so, we not only contribute to the academic and practical understanding of data contamination issues but also highlight the pressing need for strategies that mitigate these risks.
@@
  • We categorize contamination detection into two broad types of use cases vastly differing in the techniques involved, namely open-data and closed-data contamination detection, and review all existing work in each category. We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
  • We introduce LLMSanitize, an open-source Python library aimed at helping the community centralize a shared implementation of all contamination detection methods.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to open-data contamination detection. Section 3 addresses the issue of closed-data contamination detection, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices. Section 5 briefly introduces our library LLMSanitize implementing contamination detection algorithms, which is further described in Appendix.
2 Open-Data Contamination Detection
In the open-data setup, we assume that the (pre)-training set is known. Given a training dataset $D$ and an evaluation dataset $D_{E}$, the objective is to check whether $D\cap D_{E}$ is empty. In the following, we review existing open-data contamination detection methods, which directly compare ensembles of data points.
@@ 2.1 String Matching
The most straightforward method to compute overlap between textual datasets is string matching or analyzing the set of words intersecting two texts.
Several LLM pre-training reports assess contamination levels by using string matching techniques. GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. An average of 3.2% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training is discovered. GPT-3 (Brown et al., 2020) scans data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). Some Wikipedia language modeling benchmarks, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) are nearly entirely contaminated. PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Both GPT-3 and PaLM authors conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees. GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data by computing 50-characters substring collisions. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results. Llama-2 (Touvron et al., 2023) refines the methods above by using tokens instead of words to measure contamination, and asserts that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. From there, the contamination level of a sample is defined as its percentage of contaminated tokens.
Other studies analyze overlap between training sets and evaluation sets to assess contamination levels. Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the C4 pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%. Li (2023b) compute the METEOR (Banerjee and Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and consider those with scores over 0.75 as contaminated. Contamination levels ranging from 1% to 47% across popular benchmarks are identified. Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then splitting retrieved documents into 13-gram chunks, and computing overlap metrics between these chunks and the evaluation data point’s chunks. TruthfulQA (Lin et al., 2021) exhibits high overlap with the pre-training datasets.
@@ 2.2 Embeddings Similarity
Beyond simple string matching, computing cosine similarity between embeddings offers an attractive alternative, as this process is more robust to surface-form vocabulary changes like paraphrases.
Lee et al. (2023) prevent contamination in their Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers and Gurevych, 2019)) greater than 80% against any training item. Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their CodeExercises dataset and the evaluation set HumanEval (Chen et al., 2021). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from (Riddell et al., 2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). Widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 Model-based Open-Data Contamination Detection
Embeddings similarity is limited by its need to choose an appropriate similarity threshold. To improve robustness, some methods attempt to use LLMs themselves to detect contamination in LLMs, leveraging their high performance in paraphrasing and paraphrase detection (Witteveen and Andrews, 2019; Abaskohi et al., 2023).
@@ 2.3 Model-based Open-Data Contamination Detection
3 Closed-Data Contamination Detection
In closed-data contamination detection, the training data is unknown. Formally, a model $M$ has been trained on a training set $D_{M}$ (the pre-training set is specific to each LLM) and is applied to an evaluation dataset $D_{E}$. $D_{M}$ is not accessible, and researchers have to use tools analyzing $M$’s behavior on $D_{E}$.
@@ 3.1 Performance Analysis
We first review a line of research assessing model contamination through simple performance analysis. $M$ is applied to different evaluation datasets, and a sudden drop in performance suggests previous contamination.
The most straightforward performance analysis consists in plotting performance against time, which leverages the fact that evaluation datasets released after a LLM is pre-trained should not be contaminated. In the code domain, Roberts et al. (2023) investigate the performance of GPT-3.5 and GPT-4 on datasets released before and after the training cutoff for both models. The analysis underscores a statistically significant positive correlation for problems catalogued on GitHub prior to the GPT models’ training cut-off, suggesting that their inclusion in the training data enhanced their performance on related tasks. Concurrent to their work, Li and Flanigan (2023) also plot the performance of several LLMs including GPT-3 (Brown et al., 2020) and open-source LLMs on general natural language understanding tasks against time, specifically before and after each LLM’s training data cutoff date. They find that LLMs are much more likely to improve on a trivial majority baseline for datasets released before their cutoff date, suggesting some contamination. While evaluating on fresh data guarantees a contamination-free performance evaluation, it is hard to scale this method in time as it requires frequent update of evaluation sets. Besides, the work of Cao et al. (2024) shows that evaluating on data posterior to the LLM training cutoff time may not always show a decrease in performance: in the code domain, the opposite may happen, complicating the usage of training cutoffs as a contamination detection method.
Jiang et al. (2024b) delve deep into the impact of evaluation data leakage within the pre-training set by running a series of controlled pre-training experiments with GPT-2 (Radford et al., 2019). They pre-train from scratch GPT-2 in three different ways: with the original pre-training set, with additional input texts from four evaluation datasets (SST-2, MMLU (Hendrycks et al., 2020), CNN/DM (Hermann et al., 2015) and SQuAD-v1 (Rajpurkar et al., 2016)), and with input texts and labels from these same evaluation datasets. Contaminated pre-training clearly boosts performance on all downstream tasks. Findings also reveal a U-shaped relationship between the number of times that contaminated data points are included in the pre-training, and model performance. Authors also show that the n-gram contamination detection method from Llama-2 Touvron et al. (2023) is not effective at detecting contaminated data points, calling for more robust methods.
3.2 Model Memorization